### v1

# Molecular optomechanics with atomic antennas

Abstract
A typical surface-enhanced Raman scattering (SERS) system relies on deeply subwavelength field localization in nanoscale plasmonic cavities to enhance both the excitation and emission of Raman-active molecules. Here, we demonstrate that a germanium-vacancy (GeV) defect in diamond can efficiently mediate the excitation process, by acting as a bright atomic antenna. At low temperatures, the GeV’s low dissipation allows it to be efficiently populated by the incident field, resulting in a thousand-fold increase in the efficiency of Raman scattering. Additionally, we show that atomic antenna-enhanced Raman scattering can be distinguished from conventional SERS by tracing the dependence of Stokes intensity on input power, and the pronounced antibunching of the Raman emission.
Raman spectroscopy is a highly versatile technique for detecting and studying molecules and proteins, and despite over four decades of research, it continues to provoke new research questions and yield surprising experimental results. In particular, in the last decade a new class of plasmonic nano- and picocavities [1, 2, 3, 4] offering extreme field localisation has enabled a new level of capability. This includes imaging the internal structure of molecules with sub-nanometer resolution [5], observing the stretching of individual atomic bonds [6], and ultimately controlling their chemistry [7], while relying on only visible light. These studies have also extended beyond plasmonic systems, with several groups pursuing the realisation of hybrid cavities, which inherit strong localisation of light from plasmonics, and small dissipation from dielectrics [8, 9, 10].
Figure 1: (a) Schematic and (b) idealized spectrum of atomic antenna-enhanced Raman scattering. The incident laser at frequency $\omega_{l}$ couples to an atomic antenna, which localizes the electric field on the nearby Raman-active molecule. The Raman emission (here shown as a Stokes line) from the molecule is mediated by the plasmonic antenna.
These developments in alternative platforms have been accompanied by novel theoretical formalisms. In particular, “molecular optomechanics” [11, 12, 13] aims to describe cavity-enhanced Raman scattering by embracing a quantum-mechanical picture of both the optical mode (or modes) and the molecular vibrations. To date, molecular optomechanics has provided insights into various fundamental phenomena such as the optical spring effect and linewidth narrowing [1], dynamical instabilities [14], and correlations between the Stokes and anti-Stokes emission [12, 15]. It has also provided a convenient formalism for describing upconversion of IR photons via Raman transitions [16, 17, 18], and offered insights into the effects of molecular anharmonicities [19, 20], and coherences between vibrations of neighboring molecules [21].
The framework of molecular optomechanics also transparently describes the dual effect of plasmonic cavities in Tip-Enhanced (TERS) and Surface-Enhanced (SERS) Raman Specstroscpy systems [22, 23]. Specifically, cavities serve to (i) localize the field around the molecule, enhancing its driving to a virtual state, and (ii) boost the Stokes and anti-Stokes emission rates by modifying the density of electromagnetic states. In conventional plasmonic SERS, both effects are associated with either one, or several modes of the plasmonic cavity [24, 25], which have linewidths $\kappa$ similar or slightly smaller than the mechanical frequency of the molecular vibrations ($\kappa\lesssim\omega_{b}$). The efficiency of this process can be described by a simple figure of merit $Q/V_{\text{eff}}^{2}$, defined through the effective mode volume $V_{\text{eff}}$ and quality factor $Q$ of the optical mode.
Here we propose an alternative platform to these designs, in the form of a system incorporating an atomic defect — specifically, a germanium vacancy (GeV) in diamond — and a plasmonic nanoparticle (see schematic in Fig. 1(a)). At low temperatures, GeV defects offer two key characteristics ideal for a Raman-enabling antenna: a very low decoherence rate of order $\gamma/2\pi\approx 100$ MHz (about 4 times larger than its lifetime transformed decay rate $\gamma_{0}$), and an unprecedented degree of localisation of the electromagnetic fields. The former means that by acting as an atomic antenna [26], GeV can be far more efficiently excited than any mode of a plasmonic cavity, while the latter ensures that it can drive Raman processes in nearby molecules. At first sight, one might hope to replace both resonant functions of the plasmonic nanoparticle described above with the atomic antenna. However, the atomic linewidth is too narrow to accommodate both the pump and Raman-shifted lines, and it must be partnered with another component, e.g. a gold nanoparticle.
In this work we assess the viability of this setup, which we dub atomic antenna-enhanced Raman scattering. We ask whether atomic antennas can provide effective excitation of the Raman processes, and how is this reflected by the parameters of molecular optomechanics? Can the system offer stronger perturbation of vibrational populations, perhaps leading to dynamical instabilities? How will the efficiency of the atomic antenna-enhanced Raman scattering compare to that of conventional plasmonic SERS, and will the intrinsic nonlinearity of the atomic antenna bring about new effects that would differentiate its effects from those of a conventional plasmonic setup?
The paper is constructed as follow: in Section I we modify the formalism of molecular optomechanics to explore the differences between the atomic antenna and conventional SERS systems, and compare the two for a set of characteristic parameters in Section II. Finally, in Section III we discuss the effects of the incoherently nonlinear response of the atomic antenna and the resulting characteristics of Raman response.
I Atomic antenna-enhanced Raman scattering
I.1 Coherent dynamics
We consider a Raman dipole $\hat{\mathbf{p}}_{R}$, induced in the molecule positioned at $\mathbf{r}_{m}$ by the electric field $\hat{\mathbf{E}}_{\sigma}$ generated by the excited atomic antenna: $\hat{\mathbf{p}}_{R}=\hat{x}R:\hat{\mathbf{E}}_{\sigma}$. Here $R$ is the Raman tensor, and $\hat{x}$ is the generalized coordinate operator which describes the coordinates of the atoms in the molecule. In the lowest-order approximation [27], we quantize it as a one-dimensional harmonic oscillator, with zero-point fluctuation $Q_{0}=\sqrt{\hbar/(2\omega_{b})}$, frequency $\omega_{b}$, and bosonic operators $\hat{b}/\hat{b}^{\dagger}$ describing the annihilation/creation of the optical phonon. The Raman dipole interacts with the electric field of a particular mode of the plasmonic cavity $\hat{\mathbf{E}}_{a}$, as described by the interaction Hamiltonian [11, 12]
$\hat{H}_{I}$ $=-\frac{1}{2}\hat{\mathbf{p}}_{R}\cdot\hat{\mathbf{E}}_{a}(\mathbf{r}_{m})$
$=-\frac{1}{2}\Big{(}\hat{\mathbf{E}}_{\sigma}(\mathbf{r}_{m})\Big{)}_{i}R_{ij}\hat{x}\Big{(}\hat{\mathbf{E}}_{a}(\mathbf{r}_{m})\Big{)}_{j}$
$=-\frac{1}{2}R_{ij}Q_{0}(\hat{b}+\hat{b}^{\dagger})\Big{(}\hat{\mathbf{E}}_{\sigma}(\mathbf{r}_{m})\Big{)}_{i}\Big{(}\hat{\mathbf{E}}_{a}(\mathbf{r}_{m})\Big{)}_{j}.$ (1)
To express the electric field from the atomic antenna positioned at $\mathbf{r}_{\sigma}$, we model it as a two-level system with frequency $\omega_{\sigma}$, and pseudo-spin operators $\hat{\sigma}=|g\rangle\langle e|$. Re-interpreting the field amplitudes as being generated by the atomic transition dipole moment with matrix element $\mathbf{d}_{\sigma}=\langle e|\hat{\mathbf{d}}|g\rangle$, and using the electromagnetic Green’s function of the system $\mathbf{G}(\mathbf{r}_{\text{src}},\mathbf{r}_{\text{dest}},\omega)$, we find
[…]
The electric field operator of the plasmonic cavity mode is expressed in terms of the associated classical mode function $\mathbf{e}_{a}$ and annihilation/creation operators $\hat{a}/\hat{a}^{\dagger}$ for the plasmon polariton, such that
[…]
where $\mathbf{e}_{a}$ carries the fast harmonic time dependence, information on the spatial distribution, and the normalisation conditions.
The interaction Hamiltonian (we omit the position and frequency dependencies)
[…]
is accompanied by the free-energy Hamiltonian $\hat{H}_{0}$ terms describing the energy of the vibrations, atomic antenna and the optical cavity
$\hat{H}_{0}=\hbar\omega_{b}\hat{b}^{\dagger}\hat{b}+\hbar\omega_{\sigma}\hat{\sigma}^{\dagger}\hat{\sigma}+\hbar\omega_{a}\hat{a}^{\dagger}\hat{a},$ (5)
and a term $\hat{H}_{\text{coh}}$ describing the coherent driving of the atomic antenna by an incident laser
$\hat{H}_{\text{coh}}=\hbar\Omega\left[e^{-i\omega_{l}t}\hat{\sigma}+e^{i\omega_{l}t}\hat{\sigma}^{\dagger}\right],$ (6)
where
$\hbar\Omega=|\mathbf{d}_{\sigma}|t_{p}\sqrt{\frac{4I}{c\varepsilon_{0}}},$ (7)
is the coherent driving amplitude from the laser with intensity $I$. Note that we explicitly account for the change in the intensity of the $p$-polarized incident laser due to the transmission from air to the diamond substrate, by including the Fresnel transmission coefficient $t_{p}$ in the definition of $\Omega$.
I.2 Dissipative dynamics
To account for the dissipation in the system, we embrace the formalism of open quantum systems, in which the state of the quantum system, comprising the two-level system modeling the atomic antenna, plasmonic cavity mode, and vibrations, is described with a density matrix $\hat{\rho}$, and its dynamics is governed by a master equation. The details of this description are discussed in previous contributions on molecular, and cavity optomechanics [13, 28, 29, 30, 31]. Here we only clarify that the dissipation of the optical cavity mode is incorporated into the master equation via the Gorini–Kossakowski–Sudarshan–Lindblad (GKSL) operators $(\kappa/2)\mathcal{D}[\hat{a}]$, where $\mathcal{D}[\hat{O}]\rho=2\hat{O}\rho\hat{O}^{\dagger}-\hat{O}^{\dagger}\hat{O}\rho-\rho\hat{O}^{\dagger}\hat{O}$. Similarly, the interaction of the vibrational mode with the thermal environment — both decay and incoherent pumping — is described by terms $(\Gamma/2)(n_{b}^{\mathrm{th}}+1)\mathcal{D}[\hat{b}]$ and $(\Gamma/2)n_{b}^{\mathrm{th}}\mathcal{D}[\hat{b}^{\dagger}]$, respectively, where $n_{b}^{\mathrm{th}}$ is the thermal population of the environment at the frequency $\omega_{b}$. The decay of the atomic antenna is included as $(\gamma/2)\mathcal{D}[\hat{\sigma}]$.
I.3 Linearisation and effective optomechanical coupling
The complete Hamiltonian $\hat{H}_{0}+\hat{H}_{I}+\hat{H}_{\text{coh}}$ can be linearized by realising that in the limit of weak coherent pumping (the so-called Heitler regime) and weak optomechanical interaction, the atomic antenna ($\hat{\sigma}$) will be driven by the incident laser into a predominantly coherent state with amplitude  [^1] In the original papers on molecular optomechanics, we assumed only one plasmonic mode, and attributed the Raman-generated sidebands in the optical mode with the fluctuations in $a$ around its mean.
$\alpha_{\sigma}=\frac{i\Omega}{-(\omega_{\sigma}-\omega_{l})+i\gamma/2}\approx\left\langle\hat{\sigma}\right\rangle.$ (8)
We revisit this approximation in the last section of the paper, where we explore the Mollow limit of strong pumping of the atomic antenna.
The linearized Hamiltonian, in the rotating wave approximation, and with some elementary assumptions about the isotropic nature of the Raman tensor, takes on a simpler form (see Appendix A for derivation)
$\hat{H}_{I,\text{lin}}$ $=-\underbrace{\frac{1}{2}RQ_{0}\omega_{\sigma}^{2}\mu_{0}\mathbf{e}_{a}\cdot(\mathbf{G}\cdot\mathbf{d}_{\sigma})^{*}}_{\hbar g_{0,\sigma}}\alpha_{\sigma}\left(\hat{a}+\hat{a}^{\dagger}\right)(\hat{b}+\hat{b}^{\dagger})$
$=-\hbar g_{0,\sigma}\alpha_{\sigma}\left(\hat{a}+\hat{a}^{\dagger}\right)(\hat{b}+\hat{b}^{\dagger}).$ (9)
The coefficient $g_{\sigma}=g_{0,\sigma}\alpha_{\sigma}$ is the effective optomechanical coupling, and accounts for the population of the atomic antenna. In conventional cavity optomechanics, these populations can be very large, and compensate for the normally weak $g_{0}$ [28]. Conversely, plasmonic cavities are characterized by very large losses, and cannot be easily populated far above one plasmon ($|\alpha_{a}|=1$). For an atomic antenna, the coherence similarly cannot exceed $|\alpha_{\sigma}|=1$, since the population is by definition limited to unity.
I.4 Emission
Let us consider the spectrum of emission from the system, assuming that the inelastically scattered Stokes and anti-Stokes photons are emitted from the molecule into the plasmonic cavity. If the cavity emits light as a dipolar scatterer, its incoherent emission spectrum will be given by
$S_{a}(\omega)=\eta_{\text{out}}\omega^{4}\int_{-\infty}^{\infty}\textrm{d}\tau~{}e^{i\omega\tau}\braket{\hat{a}^{\dagger}(\tau)\hat{a}(0)}.$ (10)
We appreciate that the frequency prefactor is typically not included in the definitions of the emission spectra, but we include it here to directly relate $S_{a}(\omega)$ with the spectral density of the emitted power (see discussion in Appendix B). In Appendix C, we show that this spectrum is given by
$S_{a}$ $(\omega)=2\omega^{4}|g_{\sigma}|^{2}$
×$~{}~{}\times\Bigg{[}\frac{n_{b}}{(\Delta_{a}-\omega_{b})^{2}+(\kappa/2)^{2}}\frac{\Gamma}{(\omega_{b}+\omega_{l}-\omega)^{2}+(\Gamma/2)^{2}}$
$~{}~{}~{}~{}~{}~{}+\frac{n_{b}+1}{(\Delta_{a}+\omega_{b})^{2}+(\kappa/2)^{2}}\frac{\Gamma}{(\omega_{b}-\omega_{l}+\omega)^{2}+(\Gamma/2)^{2}}\Bigg{]},$ (11)
where $\Delta_{a}=\omega_{a}-\omega_{l}$, and $n_{b}=\langle\hat{b}^{\dagger}\hat{b}\rangle$ is the population of molecular vibrations, discussed below. At the Stokes ($\omega_{S}=\omega_{l}-\omega_{b}$) and anti-Stokes frequencies ($\omega_{aS}=\omega_{l}+\omega_{b}$) the second, and first terms dominate, respectively:
$S_{a}(\omega_{S})$ $\approx\frac{2\omega_{S}^{4}|g_{\sigma}|^{2}(n_{b}+1)}{[(\Delta_{a}+\omega_{b})^{2}+(\kappa/2)^{2}]\Gamma},$ (12)
$S_{a}(\omega_{aS})$ $\approx\frac{2\omega_{aS}^{4}|g_{\sigma}|^{2}n_{b}}{[(\Delta_{a}-\omega_{b})^{2}+(\kappa/2)^{2}]\Gamma}.$ (13)
I.5 Vibrational populations
In the absence of optical driving, the population of vibrations is determined by the thermal environment, leading to $n_{b}=n_{b}^{\text{th}}$. However, the Stokes and anti-Stokes transitions continuously add, and remove phonons from the molecule, respectively. The rates of these processes ($\Gamma_{\pm}$) are proportional to the emission intensities $S_{a}(\omega_{S})$ and $S_{a}(\omega_{aS})$ respectively, and can be calculated as [33, 28]
$\Gamma_{\pm}=\frac{g_{\sigma}^{2}\kappa}{(\Delta_{a}\pm\omega_{b})^{2}+(\kappa/2)^{2}}.$ (14)
We can then formulate the rate equation for the phonon population as (see derivation in Ref. 13)
$\frac{\textrm{d}}{\textrm{d}t}n_{b}=-(\Gamma+\Gamma_{-})n_{b}+\Gamma_{+}(n_{b}+1)+\Gamma n_{b}^{\text{th}},$ (15)
with the steady-state solution
$n_{b,ss}=\frac{\Gamma n_{b}^{\text{th}}}{\Gamma+\Gamma_{-}-\Gamma_{+}}+\frac{\Gamma_{+}}{\Gamma+\Gamma_{-}-\Gamma_{+}}.$ (16)
The effective dissipation rate of the phonon mode is also modified by these processes as $\Gamma\rightarrow\Gamma_{\text{eff}}=\Gamma+\Gamma_{-}-\Gamma_{+}$. Both effects (the changes in the steady-state phonon population and the linewidth) should be included into (Eq. (I.4)) for the spectrum by substituting $n_{b}\rightarrow n_{b,ss}$ and $\Gamma\rightarrow\Gamma_{\text{eff}}$.
Typical Raman scattering experiments are carried out at room temperature, where the thermal phonon populations $n_{b}^{\text{th}}$ are of the order of $10^{-1}$, and at laser intensities where the modification of the dissipation rate is negligible. In fact, conclusive observation of the change of phonon population due to vibrational pumping (described by the second term on the right-hand side of Eq. (16)), was only reported in [34], by observing a quadratic dependence of the anti-Stokes intensity on pump laser power. Furthermore, observation of a build-up of phonon populations due to the optomechanical heating, represented by the reduction of the effective phonon damping rate $\Gamma_{\text{eff}}$, requires the use of pulsed lasers with large peak powers, to avoid destroying the plasmonic cavities [14].
Atomic defects, like the GeV discussed here, require low temperatures (optimally around $50$ K [26]) to suppress the decoherence, and operate as high-$Q$ antennas. At those temperatures the thermal phonon population $n_{b}^{\text{th}}$ is further reduced. To estimate how strong that mechanisms can be, in the next section we estimate the parameters of a realistic scheme for atomic-antenna enhanced Raman scattering.
II Comparison to a conventional plasmonic SERS
Having developed the general framework for describing atomic antenna-enhanced Raman scattering, we now seek to compare its performance to that of a conventional SERS system, for which the formalism of molecular optomechanics was originally developed [11, 12]. For completeness, we list the expressions for relevant parameters in the latter setup in Appendix E.
In Table 1 we present a comparison of parameters characterising an atomic antenna-enhanced Raman scattering scheme (in a setup schematically represented in Fig. 1), and a conventional plasmonic SERS setup with the same nominal geometry, but without the atomic antenna, where the plasmonic cavity mode mediates both the excitation of the Raman dipole, and its emission. For simplicity, both systems use the same properties of the plasmonic cavity mode (frequency, linewidth and effective mode volume). We note that these parameters are not intended to describe a particular specific plasmonic system, but are characteristic of various nanoparticle-on-mirror and nanoparticle-in-grove platforms [29]. To obtain simple expressions, we also invoke substantial approximations, most significantly (i) assuming that the Green’s function of the system can be approximated by the free-space Green’s function, (ii) taking the coherent driving amplitude $\Omega$ as defined in Eq. (7), without accounting for additional focusing due to the nanoparticle, and (iii) treating the decay rate of the atomic antenna $\gamma$ as independent of the presence of the nanoparticle. We revisit all these approximations in Subsection II.1.
Parameters of the atomic antenna, including the transition dipole moment, are directly taken, or inferred, from Ref. [26] (see Appendix D for the detailed derivation). We assume a shallow location of the defect, 2 nm below the surface of the diamond. We also emphasize that while this work focuses on coherent color centers in diamond, optically coherent defects are observed in several systems, including molecules, rare earth dopants, nanodiamonds, and atomically-thin layered materials such as hexagonal boron nitride [35, 36, 37, 38, 39]. These systems are all suitable candidates for atomic-antenna enhanced Raman spectroscopy.
Conventional plasmonic system Atomic antenna
Parameters $(\omega_{a},\kappa)/2\pi=(498,50)$ THz,
$V_{\text{eff}}=10^{3}~{}\text{nm}^{3}\approx\lambda^{3}/(2\times 10^{5})$,
$\sigma_{\text{ext}}=10^{-13}~{}\text{m}^{2}$ $(\omega_{\sigma},\gamma_{0})/2\pi=(498~{}\text{THz},25~{}\text{MHz})$
$\gamma=4\gamma_{0}$ (at 50 K [26])
$|\mathbf{d}_{\sigma}|\approx 6.7$ D (Appendix D)
$|\mathbf{r}_{\sigma}-\mathbf{r}_{m}|=3$ nm
$RQ^{0}\approx 4\pi\varepsilon_{0}2\times 10^{-30}~{}\text{m}^{3}$$\omega_{b}/2\pi=40$ THz, $\Gamma/2\pi=1$ THz
Electric field $e_{a}\approx 1.4\times 10^{8}$ V/m (Eq. (50)) $e_{\sigma}=\omega_{\sigma}^{2}\mu_{0}G(\mathbf{r}_{\sigma},\mathbf{r}_{m})|\mathbf{d}_{\sigma}|\approx 7.5\times 10^{6}$ V/m
(1 plasmon/exciton)
Optomechanical coupling $\hbar g_{0,a}\approx 1.4\times 10^{-5}$ eV (Eq. (51)) $\hbar g_{0,\sigma}\approx 7\times 10^{-7}$ eV (Eq. (I.3))
$g_{0,a}/2\pi\approx 3.3\times 10^{9}$ Hz $g_{0,\sigma}/2\pi\approx 1.8\times 10^{8}$ Hz
Laser intensity required $I\approx 10^{3}~{}\upmu\text{W}\cdot\upmu\text{m}^{-2}$ $I\approx 6.9\times 10^{-3}~{}\upmu\text{W}\cdot\upmu\text{m}^{-2}$ ($|\alpha_{\sigma}|=1$ in
for 1 plasmon/exciton ($|\alpha_{a}|=1$ in Eq. (48)) in Eq. (8); $60$ deg incidence, $t_{p}=0.46$)
Table 1: Parameters of the systems used to compare the performance of the conventional plasmonic system and atomic antenna. The properties of the plasmonic antenna are shared between the two models, although in the case of the latter system the role of the plasmonic cavity is limited to controlling the Raman emission. Details on the properties of the atomic antenna and the addressed Raman activity are given in Appendix D; equations used to analyse the response of a conventional plasmonic system are given in Appendix E.
The laser intensity which would be required to excite a single coherent plasmon (i.e. to induce a coherent population $|\langle\hat{a}\rangle|=1$) is about $10^{2}~{}\upmu\text{W}\cdot\upmu\text{m}^{-2}$ (see discussion in Appendix E), similar to the maximum CW intensities reported in Ref. [14]. For comparison, the atomic antenna becomes saturated for laser intensities about five orders of magnitude smaller (we discuss the effects of saturation of the atomic dipolar transition in the antenna in Section III). These estimates are consistent with the reported values of power required to saturate other group-IV centers in diamond, e.g. SnV [40].
The single-photon optomechanical coupling in conventional SERS ($g_{0,a}$) is about $20$ times larger than for the atomic antenna ($g_{0,\sigma}$), and both are far smaller than the record values estimated for plasmonic picocavities [2].
This contrast of laser intensities required for driving the two systems, and the similarity of the single-photon optomechanical coupling, is the first key message of our work. It demonstrates that atomic antenna-enhanced Raman scattering is a viable alternative for conventional SERS, in particular in systems in which the use of high-intensity lasers is undesirable due to fragility of the sample, for instance to photo-induced bleaching or charge instability, or the focusing element.
We can also estimate the efficiency of down-conversion of the incident to emitted Stokes photons  [^2] A similar question of the conversion efficiency of IR to visible photon was considered by Roelli and co-authors in Ref. [16].. We derive the complete expression for this efficiency in Appendix B, and find that it can be expressed as
$\eta_{\sigma}$ $=\eta^{\prime}\frac{\left|g_{0,\sigma}\alpha_{\sigma}\right|^{2}}{I},$ (17)
where $\eta^{\prime}$ is a prefactor given in Eq. (B). Similarly, we quantify the conversion efficiency for the conventional plasmonic Raman system
$\eta_{a}=$ $\eta^{\prime}\frac{\left|g_{0,a}\alpha_{a}\right|^{2}}{I}.$ (18)
It is clear from Eqs. (17) and (18), that the difference in efficiencies between the two systems is dictated by the effective optomechanical couplings $g_{0,\sigma}\alpha_{\sigma}$ and $g_{0,a}\alpha_{a}$, normalized by the intensity of the incident laser. Exciting plasmons requires much larger laser intensities than exciting the atomic antennas (see Table 1), meaning that per intensity $|\alpha_{a}|^{2}/I\ll|\alpha_{\sigma}|^{2}/I$; on the other hand, conventional plasmonic cavities yield larger single-photon optomechanical couplings $|g_{0,a}|\gg|g_{0,\sigma}|$.
Therefore to make meaningful assessments about the resulting contrast in efficiencies, we need to perform a more careful analysis, investigating the effects of the three approximations discussed above.
II.1 Modification of the properties of atomic antenna by a nanoparticle
In the list of parameters given in Table 1, we have neglected the effects of the nanoparticle on the characteristics of the atomic antenna. Here we carry out a more quantitative discussion of these three effects, and their impact on the overall Raman scattering efficiency, supported by numerical solutions implemented in COMSOL Multiphysics software. Details of the numerical model are given in Appendix F. We stress that these calculations are only aimed at establishing the relative corrections to the driving amplitude, decay rate, and field of the atomic antenna, rather than re-assessing all the parameters given in Table 1. Therefore, we keep the remaining values, most critically the frequency and mode volume of the cavity, as constant in the following discussion, meaning that the efficiency of the conventional plasmonic SERS setup remains unchanged.
The first two effects are commonly discussed in the context of plasmon-enhanced or quenched fluorescence [42, 43]:
  • Enhancement of the coherent driving amplitude
    The plasmonic nanoparticle will modify the intensity of the incident laser at the site of the atomic antenna by a factor $K_{\text{inc}}^{2}$, enhancing its driving amplitude, and lowering the threshold for the excitation of the atomic antenna. To quantify this effect, we calculate the field distribution for illumination with a $p$-polarized plane wave incident at an angle of $60^{\circ}$, for a range of substrate-nanoparticle distances. We plot the field intensity normalized by the field intensity calculated in the absence of the nanoparticle (which already accounts for the imperfect transmission of the incident light into the substrate) in Fig. 2(a) with a solid blue line.
  • Modification of the atomic antenna dissipation via the Purcell effect
    The plasmonic nanoparticle will also modify the intrinsic dissipation pathways for the atomic antenna, both by enhancing its coupling to the radiation modes, and by introducing nonradiative decay channels via absorption in the metal. The resulting decay rate of the atomic antenna will be increased, meaning that a larger laser intensity is required for its excitation. We model this effect by calculating the scattering from a dipolar electric source embedded in the diamond substrate, and oriented vertically, and quantify the enhancement of the decay rate by calculating the sum of the radiated and absorbed power ($P_{\text{rad}}+P_{\text{abs}}$), normalized by the emitted power in the absence of the nanoparticle $P_{0}$. Since this enhancement will only modify the rate of the radiative decay of the emitter $\gamma_{0}$, the overall decay rate can be estimated as
    $\gamma\rightarrow\gamma_{\text{tot}}$ $=\underbrace{\frac{P_{\text{rad}}+P_{\text{abs}}}{P_{0}}\gamma_{0}}_{\text{radiative decay}}\,\,+\underbrace{3\gamma_{0}}_{\text{non-radiative decay}}$
    $=\left(\frac{1}{4}\frac{P_{\text{rad}}+P_{\text{abs}}}{P_{0}}+\frac{3}{4}\right)\gamma,$ (19)
    where we have again assumed that at the low temperature of 50 K, and in the absence of the nanoparticle, $\gamma=4\gamma_{0}$ [26]. We plot the resulting shortening of the antenna’s lifetime $(\gamma_{\text{tot}}/\gamma)^{-1}$ in Fig. 2(a) with a solid orange line  [^3] We note the limiting case of the decay rate enhancement from a buried electric dipole was considered in Ref. 50, where the enhancement in a similar system, with a plasmonic nanoparticle positioned on the dielectric substrate, exceeded $10^{3}$..
  • Modification of the field at the position of the Raman-active molecule
    Finally, the presence of the nanoparticle will modify the field from the atomic antenna, experienced by the molecule. To quantify this effect, we calculate the electric field from the electric dipolar source positioned 2 nm below the surface, at the distance of 1 nm above the surface of the substrate, in the gap below the nanoparticle, and normalize it by the same quantity in the absence of the nanoparticle, to find the field enhancement factor $K_{\text{mol}}$ as a function of the substrate-nanoparticle spacing $d$. We plot the result in Fig. 2(a) with a dashed green line.
The above effects enter the conversion efficiency $\eta_{\sigma}$ by (i) boosting the effective coherent population $|\alpha_{\sigma}|^{2}$ by a factor given by the field intensity enhancement $K_{\text{inc}}^{2}$, (ii) increasing the decay rate $\gamma\rightarrow\gamma_{\text{tot}}$ and thus lowering $|\alpha_{\sigma}|^{2}$ by $\gamma/\gamma_{\text{tot}}$, and (iii) modifying the optomechanical coupling coefficient $g_{0,\sigma}$ by $K_{\text{mol}}$. In Fig. 2(b) we plot the resulting corrected conversion efficiency for the atomic antenna-enhanced Raman scattering,
$\eta_{\sigma}\rightarrow\eta^{\prime}\frac{\left|g_{0,\sigma}\alpha_{\sigma}\right|^{2}}{I}K_{\text{inc}}^{2}K_{\text{mol}}^{2}\left(\frac{\gamma}{\gamma_{\text{tot}}}\right)^{2},$ (20)
and the conventional plasmonic system with parameters given in Table 1. We find that while the nanoparticle can modify the efficiency of the atomic antenna-enhanced Raman scattering via the three effects we investigated, this overall change in efficiency shown in Fig, 2(b) is fairly small.
Figure 2: Effects of the nanoparticle (NP) on the dynamics of the atomic antenna system. In (a) we plot the enhancements of the intensity of: incident field $K_{\text{inc}}^{2}$ at the antenna position (solid blue line), field from the antenna at the molecules position (dashed green line), and the reduction of the lifetime $\gamma/\gamma_{\text{tot}}$ (solid orange line) of the antenna dipole. In (b) we plot the corrected conversion efficiencies for the atomic antenna (solid red line; Eq. (20)) and conventional plasmonic system (solid gray line; Eq. (18)).
For the entire range of the substrate-nanoparticle spacings, the adjusted efficiencies of these systems are about three orders of magnitude larger than for the conventional plasmonic systems. This implies that per input photon, atomic antennas would yield about a thousand-fold increase in upconversion probability over the conventional plasmonic SERS. This dramatic contrast is the second message of our work.
III Nonlinearities in atomic antenna-enhanced Raman scattering
The results discussed in the previous section were obtained within the linearized optomechanical picture, meaning that we assumed moderate laser intensities topping up to about $7\times 10^{-3}~{}\upmu\text{W}\cdot\upmu\text{m}^{-2}$ (see Table 1), at which the coherent amplitude of the atomic antenna peaks. Here we ask what happens beyond those intensities: does the simple linearisation scheme break down? Does the atomic antenna-enhanced Stokes emission exhibit a different dependence on laser intensity than a conventional plasmonic SERS setup?
Recall that in Subsection I.1, we linearized the optomechanical Hamiltonian by replacing the pseudospin operators with their expectation values. This approximation works well in the Heitler regime, where the coherent driving amplitude $\Omega$ of the atomic antenna is far smaller than its total decay rate $\gamma$. To better understand this effect, we used QuTip [45, 46] to model the dynamics of the system described by the full (un-linearized) optomechanical Hamiltonian ($\hat{H}_{0}+\hat{H}_{\text{coh}}+\hat{H}_{I}$) and the dissipative terms described in Subsection I.2. Given how weak are the effects discussed in Section II.1, for simplicity we use the uncorrected values of the parameters listed in Table 1.
In Fig. 3(a), we plot the calculated steady-state coherence ($|\langle\hat{\sigma}\rangle|^{2}$; solid blue line) and population ($\langle\hat{\sigma}^{\dagger}\hat{\sigma}\rangle$; dashed orange line) of the antenna, and find that for weak laser power $\langle\hat{\sigma}^{\dagger}\hat{\sigma}\rangle\approx|\langle\hat{\sigma}\rangle|^{2}$. As we increase the laser power towards the Mollow regime, the atomic antenna is driven into a maximally mixed state with $\left\langle\hat{\sigma}^{\dagger}\hat{\sigma}\right\rangle=0.5$. Furthermore, the incoherent emission spectrum of the antenna
$S_{\sigma}(\omega)=\omega^{4}\int_{-\infty}^{\infty}\mathrm{d}\tau e^{i\omega\tau}\langle\hat{\sigma}^{\dagger}(\tau)\hat{\sigma}(0)\rangle,$ (21)
develops the characteristic three-peak structure of a Mollow triplet (see Fig. 3(b)).
In that regime the system cannot absorb more energy, and we expect that the intensity of the Raman processes should saturate. Indeed, the Stokes (solid green line in Fig. 3(c)) and anti-Stokes (dashed red line) intensities, calculated from the incoherent spectrum of the plasmonic cavity (Eq. (I.4)), exhibits an initial linear dependence on $I$, followed by saturation.
Figure 3: Analysis of the (a) state and (b) emission from the atomic antenna, and (c) Stokes (solid green line) and anti-Stokes (dashed red line) intensities, as a function of intensity of the pump laser. Additionally, we plot the intensity of the Stokes emission in a conventional SERS setup (solid blue line). In (a) we observe a clear saturation of the low-intensity linear behaviour when the coherent pumping rate $\Omega$ approaches the total decay rate of the atomic defect $\gamma$, lowering the coherence $|\langle\hat{\sigma}\rangle|^{2}$ (blue) and saturating the population $\langle\hat{\sigma}^{\dagger}\hat{\sigma}\rangle$ (orange), as well as the Stokes emission intensity (c). Around this point the emission spectrum of the atom develops Mollow sidebands traced with the gray dashed lines. These results have been calculated using QuTiP, with parameters matching those listed in Table 1.
In panel Fig. 3(c) we also plot, with solid blue line, the intensity of the Stokes emission from a conventional SERS setup, and find that below $I\sim 10^{-3}~{}\upmu\text{W}/\upmu\text{m}^{2}$ (the point of saturation of the atomic antenna-enhanced Stokes), this intensity is about three orders of magnitude weaker, consistent with the efficiency prediction shown in Fig. 2(b).
We note that even for the largest pump intensity for which the linearization approximation would be justified (see the value in Table 1), the phonon populations do not diverge from the thermal equilibrium values ($n_{b}^{\mathrm{th}}\approx 5\times 10^{-2}$), and the optomechanical pumping and decay rates $\Gamma_{+}$ and $\Gamma_{-}$ are several orders of magnitude smaller than $\Gamma$ ($\Gamma_{+}/\Gamma\sim 10^{-9}\gg\Gamma_{-}/\Gamma$), suggesting that we would observe neither the vibrational pumping, nor instability.
Finally, we ask whether the nonclassical statistics of the electromagnetic field generated by the atomic antenna [47, 48] — in either Heitler or Mollow regimes — yields a similarly nonclassical statistics of the emitted Raman spectrum. We expect that should be the case, as the two effects that could normally suppress the sub-Poissonian statistics — incoherent pumping, and filtering by a spectrally narrower cavity — are absent. We have confirmed this by numerically calculating the intensity correlation $g^{(2)}(0)$ of the state of the plasmonic cavity mode $\hat{a}$, which mediates the scattering. We also performed calculations of the frequency-resolved intensity correlations between the Stokes lines, and Stokes-anti-Stokes pairs [15, 49]. These results are not shown here, as in each case we find perfect antibunching of the frequency-filtered emission.
IV Summary and outlook
Thanks to a large scattering cross section and very large quality factors, atomic antennas can be very efficiently populated at relatively low laser intensities. Achieving comparable populations of a plasmonic cavity mode would require at least five orders of magnitude more intense illumination. This large intensity requirement limits the application of SERS to spectroscopy of molecules which are more sensitive to photo-dissociation/photo-bleaching, or studies requiring low temperatures. Additionally, large laser power can induce melting, or reshaping of nanoparticles.
Here we showed that the ability to efficiently excited atomic antennas means they are a promising alternative platform for low-intensity Raman scattering, although they do offer slightly smaller values of the optomechanical coupling compared to the conventional plasmonic systems. The tradeoff between these effects is captured by the efficiency of Stokes emission, which estimates the ratio of input photons which are up- and down-converted through Raman processes, and is about three orders of magnitude larger for the atomic antenna-enhanced Raman systems.
The lower intensities required for the operation of that system will also likely reduce the noise associated with nonlinear effects in plasmonic substrates.
We identify two clear signatures of the atomic-antenna enhanced Raman scattering, that can be targeted for experimental verification: the first is the dependence of the intensity of Stokes emission on the laser intensity $I$, which illustrates a clear saturation, and a significant deviation from the conventionally observed in SERS spontaneous ($\propto I$), or vibrationally pumped ($\propto I^{2}$) emission. The second is the nonclassical statistics of the emission, due to the atomic antenna transducing the incident coherent light into a stream of antibunched photons. Additionally, the sub-GHz narrowband response of the atomic antenna can be utilized to identify and filter atomic antenna-enhanced Raman scattering. By comparing photoluminescence with on- and off- resonant laser excitation of the atomic antenna, background signals can be accurately identified, similar to observations of Förster resonant excitation transfer [26].
Acknowledgements
Authors acknowledge fruitful discussions with Javier Aizpurua. MKS acknowledges funding from the Australian Research Council Discovery Early Career Researcher Award (DE220101272). AAH acknowledges funding from Q-NEXT, supported by the US Department of Energy, Office of Science, National Quantum Information Science Research Centers.
Appendix A Derivation of the interaction Hamiltonian
The linearized optomechanical Hamiltonian reads
$\hat{H}_{I,\text{lin}}=-$ $\frac{1}{2}R_{ij}Q^{0}\left(\hat{b}+\hat{b}^{\dagger}\right)$
[…]
×$\times\left[(\mathbf{e}_{a})_{j}\hat{a}+(\mathbf{e}_{a})_{j}^{*}\hat{a}^{\dagger}\right].$ (22)
In the rotating wave approximation (RWA), we drop the co-rotating terms like $\hat{a}\alpha_{\sigma}$ and $\hat{a}^{\dagger}\alpha_{\sigma}^{*}$, and write
$\hat{H}_{I,\text{lin}}$ $=-\frac{1}{2}R_{ij}Q^{0}\omega_{\sigma}^{2}\mu_{0}$
×$\times\left[\hat{a}^{\dagger}\alpha_{\sigma}(\mathbf{e}_{a})_{j}(\mathbf{G}(\mathbf{r}_{\sigma},\mathbf{r}_{m},\omega_{\sigma})\cdot\mathbf{d}_{\sigma})^{*}_{i}+\text{h.c.}\right]\left(\hat{b}+\hat{b}^{\dagger}\right).$ (23)
Taking the simplest model for an isotropic Raman tensor $R_{ij}=R\delta_{ij}$ we find
[…]
×$~{}~{}~{}\times\left(\hat{b}+\hat{b}^{\dagger}\right).$ (24)
We can choose the phase of the incident laser, or $\alpha_{\sigma}$, in such a way that $\alpha_{\sigma}\mathbf{e}_{a}\cdot(\mathbf{G}\mathbf{d}_{\sigma})^{*}$ becomes a real number, arriving at the interaction Hamiltonian given in Eq. (I.3).
Appendix B Power, flux and efficiency of Raman scattering
In the main text, we introduced the expressions for the emission spectra as given by the Fourier transform of a two-time correlator, multipled by $\omega^{4}$. This allowed us to trace the general trends of the emission intensities. However, to learn about the efficiencies of the Raman processes, here we carry out a more careful analysis.
We start by calculating the power carried via the Stokes emission. To this end, we adapt the expresssion for the total power emitted from a two-level system via a dipolar electric transition $\mathbf{d}_{\sigma}$ [51]:
$P=$ $2\varepsilon_{0}c\int\text{d}\phi\int\text{d}\theta\sin(\theta)r^{2}\int\text{d}\omega\frac{f(r,\theta,\omega)}{2\pi}$
×$\times\int\text{d}\tau e^{i\omega\tau}\langle\hat{\sigma}^{\dagger}(0)\hat{\sigma}(\tau)\rangle$ (25)
where
$f(r,\theta,\omega)=\left(\frac{\omega^{2}|\mathbf{d}_{\sigma}|}{4\pi\varepsilon_{0}c^{2}}\frac{\sin(\theta)}{r}\right)^{2},$ (26)
carries information about the scattering patern expressed in spherical coordinates $(r,\theta,\phi)$, assuming that the dipole is oriented along the $\hat{z}$ axis. With the plasmonic cavity mode emitting as an electric dipole $d_{a}=|\mathbf{d}_{a}|$, and described using operators $\hat{a}$, we rewrite these expressions as
$P=$ $2\varepsilon_{0}c\int\text{d}\phi\int\text{d}\theta\sin(\theta)r^{2}\int\text{d}\omega\frac{f(r,\theta,\omega)}{2\pi}$
×$\times\int\text{d}\tau e^{i\omega\tau}\langle\hat{a}^{\dagger}(0)\hat{a}(\tau)\rangle$ (27)
and
$f(r,\theta,\omega)=\left(\frac{\omega^{2}|{d}_{a}|}{4\pi\varepsilon_{0}c^{2}}\frac{\sin(\theta)}{r}\right)^{2}.$ (28)
We now integrate the frequencies around the Stokes emission frequency $\omega_{S}$, using Eq. (I.4) to turn the last part turns of the expression for the power to
$\int\text{d}\omega\frac{f(r,\theta,\omega)}{2\pi}\int\text{d}\tau e^{i\omega\tau}\langle\hat{a}^{\dagger}(0)\hat{a}(\tau)\rangle$
$\approx\frac{f(r,\theta,\omega_{S})}{2\pi}\int\text{d}\omega\int\text{d}\tau e^{i\omega\tau}\langle\hat{a}^{\dagger}(0)\hat{a}(\tau)\rangle$
$\approx f(r,\theta,\omega_{S})2|g_{\sigma}|^{2}\frac{n_{b}+1}{(\Delta_{a}+\omega_{b})^{2}+(\kappa/2)^{2}},$ (29)
and the power of the Stokes emission, due to the atomic antenna-enhanced Raman scattering, to
$P_{\sigma}=$ $4\varepsilon_{0}c|g_{\sigma}|^{2}\frac{n_{b}+1}{(\Delta_{a}+\omega_{b})^{2}+(\kappa/2)^{2}}$
×$\times\int\text{d}\theta\sin(\theta)r^{2}f(r,\theta,\omega_{S})\int\text{d}\phi$
$=$ $\frac{\omega_{S}^{4}d_{a}^{2}}{4\pi^{2}\varepsilon_{0}c^{3}}|g_{\sigma}|^{2}\frac{n_{b}+1}{(\Delta_{a}+\omega_{b})^{2}+(\kappa/2)^{2}}$
×$\times\int\text{d}\theta\sin^{3}(\theta)\int\text{d}\phi.$ (30)
The flux of Stokes photons is calculated as $P_{\sigma}/(\hbar\omega_{S})$.
We can compare it to the flux of pump photons $I\sigma_{\text{spot}}/(\hbar\omega_{l})$, where $\sigma_{\text{spot}}$ denotes the cross-section of illuminating beam, and identify the efficiency of down-conversion as
$\eta_{\sigma}$ $=\frac{P_{\sigma}/(\hbar\omega_{S})}{I\sigma_{\text{spot}}/(\hbar\omega_{l})}$
$=\frac{\omega_{l}\omega_{S}^{3}d_{a}^{2}}{4\pi^{2}\varepsilon_{0}c^{3}}\frac{|g_{\sigma}|^{2}}{I\sigma_{\text{spot}}}\frac{n_{b}+1}{(\Delta_{a}+\omega_{b})^{2}+(\kappa/2)^{2}}$
×$~{}~{}~{}~{}\times\int\text{d}\theta\sin^{3}(\theta)\int\text{d}\phi.$ (31)
If the vibrational populations remain unchanged (as is the case of an atomic antenna-enhanced Raman scattering, or SERS below the vibrational pumping regime), this efficiency can be rewritten as
$\eta_{\sigma}=\eta^{\prime}\frac{g_{0,\sigma}^{2}|\langle\hat{\sigma}\rangle|^{2}}{I},$ (32)
where
$\eta^{\prime}=$ $\frac{\omega_{l}\omega_{S}^{3}d_{a}^{2}}{4\pi^{2}\varepsilon_{0}c^{3}}\frac{1}{\sigma_{\text{spot}}}\frac{n_{b}+1}{(\Delta_{a}+\omega_{b})^{2}+(\kappa/2)^{2}}$
×$\times\int\text{d}\theta\sin^{3}(\theta)\int\text{d}\phi.$ (33)
To find the value of this parameter, we use the parameters from Table 1, and set $\sigma_{\text{spot}}=10^{-12}~{}\text{m}^{2}$. The dipolar moment $d_{a}$ is estimated by equating the coherent driving amplitude $\hbar\Omega$ given in Eq. (47) with the expression for the coupling between incident laser field $\mathbf{E}_{\text{inc}}$ with a dipolar scatterer $\mathbf{d}_{a}\cdot\mathbf{E}_{\text{inc}}$. We get can estimate $\eta^{\prime}\approx 1.3\times 10^{-22}$ kg/s.
We can also derive an almost identical expression for the conventional plasmonic system, which will take the form
$\eta_{a}=\eta^{\prime}\frac{g_{0,a}^{2}|\langle\hat{a}\rangle|^{2}}{I}.$ (34)
Appendix C Derivation of the emission spectrum
To calculate the two-time correlator in Eq. (10), we consider the full Hamiltonian of the system, including the linearized interaction Hamiltonian, and free-energy terms:
$\hat{H}=\hbar\omega_{b}\hat{b}^{\dagger}\hat{b}+\hbar\underbrace{(\omega_{a}-\omega_{l})}_{\Delta_{a}}\hat{a}^{\dagger}\hat{a}-\hbar g_{\sigma}\left(\hat{a}+\hat{a}^{\dagger}\right)(\hat{b}+\hat{b}^{\dagger}).$ (35)
The Langevin equation for $\hat{a}$ reads
$\frac{\textrm{d}}{\textrm{d}t}\hat{a}(t)=\left(-i\Delta_{a}-\frac{\kappa}{2}\right)\hat{a}(t)+ig_{\sigma}[\hat{b}(t)+\hat{b}^{\dagger}(t)]+\sqrt{\kappa}\hat{a}_{\text{in}}(t),$ (36)
with the formal solution
$\hat{a}(t)$ $=ig_{\sigma}\int_{-\infty}^{t}\textrm{d}s~{}e^{\left(-i\Delta_{a}-\frac{\kappa}{2}\right)(t-s)}[\hat{b}(s)+\hat{b}^{\dagger}(s)]$
$~{}~{}~{}+\underbrace{\int_{-\infty}^{t}\textrm{d}s~{}e^{\left(-i\Delta_{a}-\frac{\kappa}{2}\right)(t-s)}\sqrt{\kappa}\hat{a}_{\text{in}}(s)}_{\hat{A}_{\text{in}}(t)}.$ (37)
Let us now remove the fast oscillation at $\omega_{b}$ from the mechanical operators by introducing $\hat{b}(s)=\hat{\tilde{b}}(s)e^{-i\omega_{b}s}$:
$\hat{a}(t)$ $=ig_{\sigma}\int_{-\infty}^{t}\textrm{d}s~{}\left[e^{\left(-i\Delta_{a}-\frac{\kappa}{2}\right)(t-s)}e^{-i\omega_{b}s}\hat{\tilde{b}}(s)+e^{\left(-i\Delta_{a}-\frac{\kappa}{2}\right)(t-s)}e^{i\omega_{b}s}\hat{\tilde{b}}^{\dagger}(s)\right]+\hat{A}_{\text{in}}(t)$
$=ig_{\sigma}\left[e^{-i\omega_{b}t}\int_{-\infty}^{t}\textrm{d}s~{}e^{\left(-i\Delta_{a}+i\omega_{b}-\frac{\kappa}{2}\right)(t-s)}\tilde{b}(s)+e^{i\omega_{b}t}\int_{-\infty}^{t}\textrm{d}s~{}e^{\left(-i\Delta_{a}-i\omega_{b}-\frac{\kappa}{2}\right)(t-s)}\tilde{b}^{\dagger}(s)\right]+\hat{A}_{\text{in}}(t)$
$\approx ig_{\sigma}\left[\frac{\hat{b}(t)}{-i\Delta_{a}+i\omega_{b}-\frac{\kappa}{2}}+\frac{\hat{b}^{\dagger}(t)}{-i\Delta_{a}-i\omega_{b}-\frac{\kappa}{2}}\right]+\hat{A}_{\text{in}}(t).$ (38)
Notably, the last step requires that the dynamics described by the integral kernel be much faster than the dynamics of $\hat{\tilde{b}}(s)$ (determined by the effective dissipation rate of the mechanical mode $\Gamma$), allowing us to approximate it as $\hat{\tilde{b}}(t)$. This Markov approximation therefore requires that $\Gamma\ll\kappa$. Another way to think about this approximation is by comparing the width of the emission Stokes/anti-Stokes lines to the spectrum of the cavity into which the molecule emits. For $\Gamma/2\pi=1$ THz, and $\kappa/2\pi=50$ THz, and the condition holds trivially.
Within the framework of the Markov approximation, the two-time correlator takes the form
$\braket{\hat{a}^{\dagger}(0)\hat{a}(s)}\approx$ $|g_{\sigma}|^{2}\left\langle\left[\frac{\hat{b}^{\dagger}(0)}{i\Delta_{a}-i\omega_{b}-\frac{\kappa}{2}}+\frac{\hat{b}(0)}{i\Delta_{a}+i\omega_{b}-\frac{\kappa}{2}}\right]\left[\frac{\hat{b}(s)}{-i\Delta_{a}+i\omega_{b}-\frac{\kappa}{2}}+\frac{\hat{b}^{\dagger}(s)}{-i\Delta_{a}-i\omega_{b}-\frac{\kappa}{2}}\right]\right\rangle$
$+\langle\hat{A}_{\text{in}}^{\dagger}(0)\hat{A}_{\text{in}}(s)\rangle$
$\approx$ $|g_{\sigma}|^{2}\left[\frac{\braket{\hat{b}^{\dagger}(0)\hat{b}(s)}}{(\Delta_{a}-\omega_{b})^{2}+(\kappa/2)^{2}}+\frac{\braket{\hat{b}(0)\hat{b}^{\dagger}(s)}}{(\Delta_{a}+\omega_{b})^{2}+(\kappa/2)^{2}}\right].$ (39)
In the above transformations, we have assumed no correlation between the input optical noise and the mechanical mode, and vanishing two-time correlators of the optical noise.
Let us now consider the simplified version of the description for the dynamics of the mechanical mode, where we assume that the vibrations are at thermal equilibrium with the environment. We therefore neglect the effects of Stokes and anti-Stokes emission which create and annihilate phonons. This effect can be easily incorporated into this picture by changing the steady-state phonon population, and effective decay rate $\Gamma$. Using the properties of the noise terms which describe the thermal environment with population $n_{b}^{\text{th}}$, that the two-time correlators are
$\braket{\hat{b}^{\dagger}(0)\hat{b}(s)}=e^{\left(-i\omega_{b}-\Gamma/2\right)s}n_{b}^{\text{th}},$ (40)
$\braket{\hat{b}(0)\hat{b}^{\dagger}(s)}=e^{\left(i\omega_{b}-\Gamma/2\right)s}(n_{b}^{\text{th}}+1).$ (41)
Since we have moved to the frame rotating with $\omega_{l}$, we need to account for it in the expression for the spectrum, which now is calculated in the lab frame. Additionally, swapping the integration $\int_{-\infty}^{\infty}\textrm{d}s\rightarrow 2\Re\int_{0}^{\infty}\textrm{d}s$, we find
$S_{a}(\omega)$ $=2\omega^{4}\Re\int_{0}^{\infty}\textrm{d}s~{}|g_{\sigma}|^{2}\left[\frac{e^{\left(-i\omega_{b}-\frac{\Gamma}{2}\right)s}n_{b}^{\text{th}}}{(\Delta_{a}-\omega_{b})^{2}+(\kappa/2)^{2}}+\frac{e^{\left(-i\omega_{b}-\frac{\Gamma}{2}\right)s}(n_{b}^{\text{th}}+1)}{(\Delta_{a}+\omega_{b})^{2}+(\kappa/2)^{2}}\right]e^{i(\omega-\omega_{l})s}$
$=2\omega^{4}\left[\frac{\eta\omega^{4}|g_{\sigma}|^{2}n_{b}}{(\Delta_{a}-\omega_{b})^{2}+(\kappa/2)^{2}}\Re\left(\frac{1}{-i\omega_{b}-i\omega_{l}+i\omega-\frac{\Gamma}{2}}\right)+\frac{|g_{\sigma}|^{2}(n_{b}+1)}{(\Delta_{a}+\omega_{b})^{2}+(\kappa/2)^{2}}\Re\left(\frac{1}{i\omega_{b}-i\omega_{l}+i\omega-\frac{\Gamma}{2}}\right)\right]$
$=2\omega^{4}|g_{\sigma}|^{2}\left[\frac{n_{b}}{(\Delta_{a}-\omega_{b})^{2}+(\kappa/2)^{2}}\frac{\Gamma}{(\omega_{b}+\omega_{l}-\omega)^{2}+(\Gamma/2)^{2}}+\frac{n_{b}+1}{(\Delta_{a}+\omega_{b})^{2}+(\kappa/2)^{2}}\frac{\Gamma}{(\omega_{b}-\omega_{l}+\omega)^{2}+(\Gamma/2)^{2}}\right].$ (42)
Appendix D Notes on the estimates for Table 1
To estimate the coherent driving amplitude and population of the atomic antenna, we need to estimate the dipole moment $\mathbf{d}_{\sigma}$ associated with the driven transition. With the transition frequency in GeV of $\omega_{\sigma}/2\pi=498$ THz (602 nm), we can estimate $|\mathbf{d}_{\sigma}|^{2}$ from the radiative contribution to the decay rate $\gamma_{0}/2\pi\approx 25$ MHz [26], which for a dipolar electric transition dipole should be given by
$\gamma_{0}=n\frac{\omega_{\sigma}^{3}|\mathbf{d}_{\sigma}|^{2}}{3\pi\varepsilon_{0}\hbar c^{3}},$ (43)
where $n$ is the refractive index of the medium ($n=2.42$ for diamond). From here we get
$|\mathbf{d}_{\sigma}|^{2}=\gamma_{0}\frac{3\pi\varepsilon_{0}\hbar c^{3}}{n\omega_{\sigma}^{3}}\approx 1.2\times 10^{-57}~{}(\text{Cm})^{2},$ (44)
or $|\mathbf{d}_{\sigma}|\approx 2.2\times 10^{-29}~{}\text{Cm}\approx 6.7$ D.
To calculate the optomechanical coupling in atomic antenna-enhanced systems, we take the free-space Green’s function, neglecting any inhomogeneities (such as interface between the dielectric and air, or off-resonant effects of metallic nanoantenna) and polarisation, and taking only its near-field component
$G_{\text{nf}}(\mathbf{r}_{\sigma},\mathbf{r}_{m})\approx\frac{1}{4\pi k^{2}|\mathbf{r}_{\sigma}-\mathbf{r}_{m}|^{3}}.$ (45)
Finally, as an example of a Raman-active molecule, we consider a $\omega_{b}/2\pi=40.5$ THz ($1351~{}\text{cm}^{-1}$) vibrational mode of the R6G [11] with Raman activity in the Gaussian units given as $3.5\times 10^{2}\varepsilon_{0}^{2}\text{\AA}^{4}\text{amu}^{-1}$, and here assumed to be equal to $R^{2}$; accounting for the transformation between Gaussian and SI units (see the SM of [11]) we find
$RQ^{0}\approx 4\pi\varepsilon_{0}2\times 10^{-30}~{}\text{m}^{3}.$ (46)
Appendix E Parameters of conventional plasmonic SERS
In this Appendix we repeat the definitions of the parameters of a conventional molecular optomechanical system, in which a single plasmonic cavity mode (with parameters denoted in the main text) mediates both the enhancement of the optical driving, and the emission from the Raman-active molecule.
  • coherent driving amplitude of the cavity mode can be estimated as
    $\Omega_{a}=\sqrt{\frac{I\sigma_{\text{ext}}}{Q4\hbar}},$ (47)
    where $Q=\omega_{a}/\kappa$ denotes the quality factor of the mode; this expression differs by a factor of $\sqrt{2}$ from the formulation given in [29]. When decoupled from (or weakly coupled to) the Raman-active molecule, and driven on resonance, the cavity settles into a coherent state with amplitude
    $\alpha_{a}=\frac{2}{\kappa}\sqrt{\frac{I\sigma_{\text{ext}}}{Q4\hbar}}.$ (48)
    Thus, reaching $|\alpha_{a}|=1$ would require laser intensity
    $I=\frac{\kappa^{2}}{4}\frac{Q4\hbar}{\sigma_{\text{ext}}}.$ (49)
  • in the approximation where the electric fields of the mode are given as
    $|\mathbf{e}_{a}|=\sqrt{\frac{\hbar\omega_{a}}{2\varepsilon_{0}V_{\text{eff}}}},$ (50)
    the optomechanical coupling takes on a simple form
    $g_{0,a}=\frac{Q^{0}R\omega_{a}}{2\varepsilon_{0}V_{\text{eff}}},$ (51)
    and the effective optomechanical coupling is defined as $g_{a}=g_{0,a}\alpha_{a}$.
  • the emission from the Raman dipole is, similarly as in the case of an atomic antenna, mediated by the plasmonic cavity mode. Therefore, the Stokes and anti-Stokes emission intensities are governed by expressions similar to those listed in Eq. (12), except for replacing the effective coupling $g_{\sigma}\rightarrow g_{a}$
    $S_{a,\text{conv}}(\omega_{S})=\frac{2\omega_{S}^{4}|g_{a}|^{2}(n_{b}+1)}{[(\Delta_{a}+\omega_{b})^{2}+(\kappa/2)^{2}]\Gamma},$ (52)
    $S_{a,\text{conv}}(\omega_{aS})=\frac{2\omega_{aS}^{4}|g_{a}|^{2}n_{b}}{[(\Delta_{a}-\omega_{b})^{2}+(\kappa/2)^{2}]\Gamma}.$ (53)
The derivations of the above expressions can be found in Refs. [11, 12, 29, 13].
Appendix F Details of the numerical simulations
To calculate the characteristics of an atomic antenna-enhanced Raman setup, we have implemented two COMSOL models which share the same geometry: a gold nanoparticle with radius of 100 nm, and permittivity $\varepsilon_{\text{Au}}=-9.53+1.51i$ ([52] at 602 nm), separated by distance $d$ from the flat diamond surface with permittivity $\varepsilon_{\text{diam.}}=n_{\text{diam.}}^{2}=5.86$. The computational domain was a sphere with radius of 3 $\mu$m, with a perfectly matching layer (PML) of 1 $\mu$m thickness. Meshing was chosen to ensure convergence, and increasingly fine towards the gap, where the element size was maxed as 0.2 nm.
We used that geometry in two settings:
  • for the calculation of $K_{\text{inc}}$ we assumed a background field associated with incident $p$-polarized planewave at 60 degree incidence, and observed the electric field at where we expect the atomic antenna to be positioned — on the axis of symmetry of the system, $2$ nm below the diamond surface,
  • for the calculation of $K_{\text{mol}}$ and $\gamma_{\text{tot}}$, an electric dipole positioned at the site of the atomic antenna, oriented vertically.